# Leveraging Computer Vision in the Intensive Care Unit (ICU) for Examining Visitation and Mobility

### v1

@@ 1 Introduction
The Intensive Care Unit (ICU) can be a traumatic experience for many patients that adversely affects their long-term physical and cognitive outcomes [1, 2]. Pain [3, 4], immobility [5, 6], and sleep deprivation [7, 8] are common sources of patient distress in the ICU [9]. Prolonged exposure to these causes can induce the onset of delirium [7, 8], post-ICU post-traumatic stress disorder (PTSD) [10, 11, 11], and increase the mortality risk [12, 13].
@@
Monitoring and assessing many aspects of patient care in the ICU is generally performed by nurses or physicians [3], which can be an error-prone practice [14, 15, 16]. Nurses are often required to work long hours and provide required shift coverage to satisfy patient demand [17]. These demands often lead to increased stress, sleep deprivation [18, 19], and instances of provider burnout [20, 21, 22]. Providing care to patients while fatigued can lead to mistakes that negatively affect patient outcomes [23, 18, 19]. Checking on the patient can also adversely affect the patient’s sleep and lead to sleep deprivation [24, 25, 26]. Patient-care interventions are frequently mentioned as a cause of poor sleep in the ICU [27]. Some studies have found that patients rarely go more than 90 minutes without receiving some form of intervention [28, 27]. Introducing AI-augmented approaches to monitor patients would reduce providers’ workload and incidents of burnout.
@@
Recent advances in AI and sensing technologies have catalyzed the development of AI-augmented solutions to monitor patient mobility and sources of sleep disruption in the critical care setting [29]. Ma et al. employed three Microsoft Kinect sensors and the Fast Region-based Convolutional Network (Fast R-CNN) to measure patient activities [15, 30]. Yeung et al. used YOLOv2 to detect and measure the duration of transitional activities, such as getting in or out of bed [31, 32]. Davoudi et al. used a collection of sensors and deep neural networks to monitor sources of sleep disruption as well as characterize the physical activity and facial expression patterns exhibited by delirious patients [33]. Using sound sensors, Davoudi et al. found that average sound pressure levels from the delirious patients’ rooms were roughly equivalent to street traffic noise. This finding is concerning because patients have reported sound as being one of the major etiologies of poor sleep quality [24].
@@
In this study, we examined how patient mobility and personnel-related disturbances correlate with measures of patients’ health, i.e., self-reported pain using DVPRS [34, 35], patient acuity [36, 37], and delirium [13]. In an IRB-approved study at the University of Florida (UF), we recruited 44 adult patients admitted to the surgical, medical, trauma, vascular, and neuromedcine ICUs. To quantify patient mobility and sleep disturbances, we employed an Azure Microsoft Kinect sensor [38] in the ICU to record privacy-preserving depth video [39] of patients during their stay. We developed and deployed a full-stack annotation tool that enabled human annotators to label the presence and posture of individuals in hospital rooms. We used the labeled data to train the YOLOv8 [40, 41] neural network to detect individuals in each depth frame and classify the activity as lying in bed or standing. After training the model on selected annotations, we ran the best-performing model on depth video frame sequences that were temporally aligned with timestamps of clinical outcomes of interest. Within each sequence, we extracted 1) the proportion of lying in bed frames, 2) the average number of people in the room, and 3) the variance of people in the room. These three metrics were examined against our clinical outcomes of interest, i.e., self-reported pain scores, patient acuity [36, 37], and delirium. Our findings revealed significant correlations between indicators of patient outcomes. Patients reported being in less pain when hospital rooms had higher visitations. In contrast, objective measurements, including patient acuity and instances of delirium, were significantly less stable when more individuals were in the room.
@@ 2.2 Data Collection
We captured depth frames in the ICU using Microsoft Kinect cameras [38], which captured 3D volumetric data of the ICU environment while safeguarding the physical identities of each person. Six compact capture and compute systems were used for data collection, each equipped with a depth camera and a compact computer. The cameras had a field of view (FOV) of $70^{\circ}$ x $60^{\circ}$ and produced 640 x 576 px image frames. Frames were saved at a frequency of 1 FPS. Clinical coordinators oversaw the deployment and use of these systems to ensure patient care was not impeded.
@@ 2.4 YOLOv8 Architecture
We used the YOLOv8 [42] architecture for our patient activity detection model. The YOLOv8’s backbone is the Darknet-53 convolutional neural network with an added C2f module [42, 43]. The C2f module, which stands for ”cross-stage partial bottleneck with two convolutions,” improves detection accuracy by combining high-level features with contextual information. To improve YOLOv8’s ability to detect smaller objects, bounding box loss is computed using the complete intersection over union (CIoU) [44] and distribution focal loss (DFL) [45] loss functions. The YOLOv8 architecture also integrates feature pyramid networks, which are adept at recognizing objects at different scales. This is beneficial for our depth frame detection task, given the varying distances of patients within the frames.
@@ 2.5 Training Pipeline
The annotated dataset in this study included 30,840 annotated depth frames from 44 patients. Model performance was evaluated using 5-fold cross-validation. Patient IDs were used to split training and test sets within each fold to prevent data leakage [46]. In each fold, 20% of the patients were allocated for testing, while the remaining patients were used for training. Experiments were conducted on a single A100 GPU. The default hyperparameters provided by Ultralytics [41] were used for our experiments. The model was trained for 100 epochs using stochastic gradient descent (SGD) and a learning rate of 0.01. We report the mean average precision (mAP), precision, recall, and F1-Score using an intersection over union threshold (IoU) of 0.5 as an average across five folds.
@@ 2.6 Characterizing mobility and sleep disturbances
The total number of detected lying in bed and standing labels within each frame was used to measure patient visitations average and variance. Walking in and out of rooms and checking different equipment can produce noise and disturb patients’ sleep. We evaluated sleep disturbances during both day and night as it has been shown that approximately 50% of ICU patients’ sleep occurs during the day [24, 27, 47].
@@ 2.7 Patient Outcomes
We assessed patient outcomes by evaluating three distinct scores, namely self-reported pain, patient acuity [36, 37], and occurrences of delirium. The patients reported pain using the Defense and Veterans Pain Rating Scale (DVPRS) [35], which is a numerical scale ranging from 0 to 10, where 0 indicated the absence of pain, 1-4 signified mild pain, 5-6 indicated moderate pain, and 7-10 represented severe pain. Patient acuity was evaluated using the methodology proposed by a recent study [36, 37]. Patient acuity was classified into two states: stable and unstable. Patients were classified as unstable if, during the specified time interval, they received continuous renal replacement therapy (CRRT), required ventilation, received a vasopressor infusion, or required a blood transfusion of 10 units or more in the past 24 hours. Delirium was evaluated using the Richmond Agitation Sedation Scale (RASS), the Confusion Assessment Method for the Intensive Care Unit (CAM-ICU), and the Glasgow Coma Scale (GCS) [48, 49].
@@ 2.8 Associating model metrics with patient Outcomes
Model metrics were recorded for each respective sequence of video frames, then evaluated for normality using D’Agostino and Pearson’s test [50, 51]. Because the null hypothesis that each sample came from a normal distribution could be rejected, the Mann-Whitney U rank test [52] was used to evaluate independence between samples. Samples were also split based on whether the record timestamps occurred during the day (7 AM - 7 PM) or night (7 PM - 7 AM). (need to mention type 1 error correction)
@@ 3.2.1 Self-Reported Pain
Patients consistently reported lower pain scores when more people were in the room (Table 3). In table 3, pain groups were combined based on whether the pain scores were classified into the no pain and mild pain groups versus the moderate and severe pain groups. The Kruskal-Wallis H-test [53] was also used to compare the four pain groups separately, but no significant difference was found ($p=0.9$). The proportion of frames with the patient lying in bed was higher for the no and mild pain group compared to the moderate and severe pain group, but the difference wasn’t statistically significant.
@@
Model Metrics
No/Mild pain Moderate/High Pain p-value
@@ 3.2.2 Patient Acuity
Model Metrics
Stable Unstable p-value
@@ 3.2.3 Delirium
Model Metrics
Non-Delirious Delirium p-value
@@ 4 Discussion
In this study, we employed the state-of-the-art YOLOv8 deep learning model [41] to detect patient postures. The results demonstrated that the model had a harder time detecting the ”lying in bed class compared to the standing class. This result aligns with our expectations because individuals lying in bed were mostly covered by thick covers that made it difficult to distinguish which features belonged to the person.
@@
Among the metrics extracted from the model, hospital room visitations average and variance exhibited substantial correlations with measures of patient outcomes. These findings are logical because noise is a crucial factor associated with sleep deprivation and the onset of delirium [54, 55], and each individual represents a unique source of noise. This observation highlights the importance of using noninvasive autonomous systems to monitor patients. Individuals who tend to wake easily during sleep may be disturbed when healthcare providers open doors or inspect equipment [24, 25, 26]. Minimizing the need for healthcare providers to enter or exit hospital rooms would allow patients to experience uninterrupted sleep and facilitate their recovery. Furthermore, implementing a noninvasive system for tracking patient mobility would allow care providers to remotely assess whether a patient is asleep and delay room entry until the patient is awake.
@@
Another intriguing discovery was the difference in the relationships identified when comparing subjective and objective measures of patient outcomes. Higher levels of hospital room visitations were linked to adverse objective measures of patient outcomes, including patient acuity and delirium. Conversely, self-reported pain exhibited a contrasting pattern. According to Boring et al., patients tend to downplay their pain when in the presence of friends or family members, as they wish to avoid causing undue worry [56]. Since we couldn’t identify anyone in each frame, it wasn’t possible to know whether the people in the room were care providers or family members. Incorporating each person’s relationship to the patient in video sequences would make it possible to explore the impact of family presence on self-reported pain records further.
@@
While a negative correlation was observed between patient outcomes and the number of people in each room, the current analysis does not provide adequate evidence to establish a causal relationship. The patient might have pressed the nurse call button [57] to request aid or their condition may have worsened to the extent that emergency medical staff were called in by the patient’s monitoring equipment. This would suggest that the number of individuals in the room resulted from the patients’ compromised outcomes. Incorporating the timing of each event that could potentially disturb a patient is needed to evaluate whether there is a causal relationship between hospital room visitations and patients’ outcomes.
@@
Using the average and variance of the number of people in an ICU room only considers one factor that can adversely affect patients’ sleep. Other stimuli, such as light or other sources of noise, have also been shown to impact patients’ sleeping habits [24, 25, 26]. Incorporating real-time data from noise and light sensors would improve our system’s capacity to identify the critical factors contributing to sleep deprivation and notify care providers of potentially serious events, such as the onset of delirium.
@@
There are advantages and disadvantages associated with the utilization of depth cameras within an ICU as opposed to RGB cameras. Depth cameras provide 3D information about a scene and preserve the privacy of each person in a scene as long as the resolution is low. However, identifying which person is the patient can be a challenging endeavor when employing a depth camera. In the future, we will incorporate patient tracking using one of the object tracking architectures that have been integrated into the YOLOv8 model [58, 59]. Using tracking information, we can follow the person lying in the hospital bed and evaluate their posture whenever they move around the hospital room. Knowing which person the patient is will enable better mobility characterization throughout the patient’s recovery period in the ICU. We can establish a comprehensive mobility index by amalgamating this data with information from other modalities, such as accelerometer data. This index would be valuable for preventing adverse events and enhancing patient outcomes.

### v2

# references.bib

@@ 1 Introduction
The Intensive Care Unit (ICU) can be a traumatic experience for many patients that adversely affects their long-term physical and cognitive outcomes [halvorsen_patients_2022, cuzco_patients_2022]. Pain [barr_clinical_2013, nordness_current_2021], immobility [adler_early_2012, tipping_effects_2017], and sleep deprivation [weinhouse_bench--bedside_2009, boyko_sleep_2012] are common sources of patient distress in the ICU [devlin_clinical_2018]. Prolonged exposure to these causes can induce the onset of delirium [weinhouse_bench--bedside_2009, boyko_sleep_2012], post-ICU post-traumatic stress disorder (PTSD) [davydow_posttraumatic_2008, righy_prevalence_2019, righy_prevalence_2019], and increase the mortality risk [pisani_days_2009, ouimet_incidence_2007].
@@
Monitoring and assessing many aspects of patient care in the ICU is generally performed by nurses or physicians [barr_clinical_2013], which can be an error-prone practice [beecroft_sleep_2008, ma_measuring_2017, wang_nurses_2010]. Nurses are often required to work long hours and provide required shift coverage to satisfy patient demand [betsiou_nursing_2022]. These demands often lead to increased stress, sleep deprivation [weaver_sleep_2016, ramadan_association_2014], and instances of provider burnout [papazian_high-level_2023, jun_relationship_2021, reader_burnout_2008]. Providing care to patients while fatigued can lead to mistakes that negatively affect patient outcomes [barker_effects_2011, weaver_sleep_2016, ramadan_association_2014]. Checking on the patient can also adversely affect the patient’s sleep and lead to sleep deprivation [al_mutair_sleep_2020, bihari_factors_2012, eliassen_sleep_2011]. Patient-care interventions are frequently mentioned as a cause of poor sleep in the ICU [beck_edvardsen_promoting_2020]. Some studies have found that patients rarely go more than 90 minutes without receiving some form of intervention [ritmala-castren_sleep_2015, beck_edvardsen_promoting_2020]. Introducing AI-augmented approaches to monitor patients would reduce providers’ workload and incidents of burnout.
@@
Recent advances in AI and sensing technologies have catalyzed the development of AI-augmented solutions to monitor patient mobility and sources of sleep disruption in the critical care setting [gholami_ai_2018]. Ma et al. employed three Microsoft Kinect sensors and the Fast Region-based Convolutional Network (Fast R-CNN) to measure patient activities [ma_measuring_2017, girshick_fast_2015]. Yeung et al. used YOLOv2 to detect and measure the duration of transitional activities, such as getting in or out of bed [yeung_computer_2019, redmon_yolo9000_2017]. Davoudi et al. used a collection of sensors and deep neural networks to monitor sources of sleep disruption as well as characterize the physical activity and facial expression patterns exhibited by delirious patients [davoudi_intelligent_2019]. Using sound sensors, Davoudi et al. found that average sound pressure levels from the delirious patients’ rooms were roughly equivalent to street traffic noise. This finding is concerning because patients have reported sound as being one of the major etiologies of poor sleep quality [al_mutair_sleep_2020].
@@
In this study, we examined how patient mobility and personnel-related disturbances correlate with measures of patients’ health, i.e., self-reported pain using DVPRS [polomano_psychometric_2016, buckenmaier_iii_preliminary_2013], patient acuity [shickel_deep_2021, ren_development_2020], and delirium [ouimet_incidence_2007]. In an IRB-approved study at the University of Florida (UF), we recruited 44 adult patients admitted to the surgical, medical, trauma, vascular, and neuromedcine ICUs. To quantify patient mobility and sleep disturbances, we employed an Azure Microsoft Kinect sensor [zhang_microsoft_2012] in the ICU to record privacy-preserving depth video [chou_privacy-preserving_2018] of patients during their stay. We developed and deployed a full-stack annotation tool that enabled human annotators to label the presence and posture of individuals in hospital rooms. We used the labeled data to train the YOLOv8 [redmon_you_2016, jocher_yolo_2023] neural network to detect individuals in each depth frame and classify the activity as lying in bed or standing. After training the model on selected annotations, we ran the best-performing model on depth video frame sequences that were temporally aligned with timestamps of clinical outcomes of interest. Within each sequence, we extracted 1) the proportion of lying in bed frames, 2) the average number of people in the room, and 3) the variance of people in the room. These three metrics were examined against our clinical outcomes of interest, i.e., self-reported pain scores, patient acuity [shickel_deep_2021, ren_development_2020], and delirium. Our findings revealed significant correlations between indicators of patient outcomes. Patients reported being in less pain when hospital rooms had higher visitations. In contrast, objective measurements, including patient acuity and instances of delirium, were significantly less stable when more individuals were in the room.
@@ 2.2 Data Collection
We captured depth frames in the ICU using Microsoft Kinect cameras [zhang_microsoft_2012], which captured 3D volumetric data of the ICU environment while safeguarding the physical identities of each person. Six compact capture and compute systems were used for data collection, each equipped with a depth camera and a compact computer. The cameras had a field of view (FOV) of $70^{\circ}$ x $60^{\circ}$ and produced 640 x 576 px image frames. Frames were saved at a frequency of 1 FPS. Clinical coordinators oversaw the deployment and use of these systems to ensure patient care was not impeded.
@@ 2.4 YOLOv8 Architecture
We used the YOLOv8 [terven_comprehensive_2023] architecture for our patient activity detection model. The YOLOv8’s backbone is the Darknet-53 convolutional neural network with an added C2f module [terven_comprehensive_2023, redmon_yolov3_2018]. The C2f module, which stands for ”cross-stage partial bottleneck with two convolutions,” improves detection accuracy by combining high-level features with contextual information. To improve YOLOv8’s ability to detect smaller objects, bounding box loss is computed using the complete intersection over union (CIoU) [zheng_distance-iou_2020] and distribution focal loss (DFL) [li_generalized_2020] loss functions. The YOLOv8 architecture also integrates feature pyramid networks, which are adept at recognizing objects at different scales. This is beneficial for our depth frame detection task, given the varying distances of patients within the frames.
@@ 2.5 Training Pipeline
The annotated dataset in this study included 30,840 annotated depth frames from 44 patients. Model performance was evaluated using 5-fold cross-validation. Patient IDs were used to split training and test sets within each fold to prevent data leakage [lekadir_future-ai_2021]. In each fold, 20% of the patients were allocated for testing, while the remaining patients were used for training. Experiments were conducted on a single A100 GPU. The default hyperparameters provided by Ultralytics [jocher_yolo_2023] were used for our experiments. The model was trained for 100 epochs using stochastic gradient descent (SGD) and a learning rate of 0.01. We report the mean average precision (mAP), precision, recall, and F1-Score using an intersection over union threshold (IoU) of 0.5 as an average across five folds.
@@ 2.6 Characterizing mobility and sleep disturbances
The total number of detected lying in bed and standing labels within each frame was used to measure patient visitations average and variance. Walking in and out of rooms and checking different equipment can produce noise and disturb patients’ sleep. We evaluated sleep disturbances during both day and night as it has been shown that approximately 50% of ICU patients’ sleep occurs during the day [al_mutair_sleep_2020, beck_edvardsen_promoting_2020, closs_patients_1992].
@@ 2.7 Patient Outcomes
We assessed patient outcomes by evaluating three distinct scores, namely self-reported pain, patient acuity [shickel_deep_2021, ren_development_2020], and occurrences of delirium. The patients reported pain using the Defense and Veterans Pain Rating Scale (DVPRS) [buckenmaier_iii_preliminary_2013], which is a numerical scale ranging from 0 to 10, where 0 indicated the absence of pain, 1-4 signified mild pain, 5-6 indicated moderate pain, and 7-10 represented severe pain. Patient acuity was evaluated using the methodology proposed by a recent study [shickel_deep_2021, ren_development_2020]. Patient acuity was classified into two states: stable and unstable. Patients were classified as unstable if, during the specified time interval, they received continuous renal replacement therapy (CRRT), required ventilation, received a vasopressor infusion, or required a blood transfusion of 10 units or more in the past 24 hours. Delirium was evaluated using the Richmond Agitation Sedation Scale (RASS), the Confusion Assessment Method for the Intensive Care Unit (CAM-ICU), and the Glasgow Coma Scale (GCS) [reznik_fluctuations_2020, jain_glasgow_2018].
@@ 2.8 Associating model metrics with patient Outcomes
Model metrics were recorded for each respective sequence of video frames, then evaluated for normality using D’Agostino and Pearson’s test [dagostino_tests_1973, dagostino_omnibus_1971]. Because the null hypothesis that each sample came from a normal distribution could be rejected, the Mann-Whitney U rank test [mann_test_1947] was used to evaluate independence between samples. Samples were also split based on whether the record timestamps occurred during the day (7 AM - 7 PM) or night (7 PM - 7 AM). (need to mention type 1 error correction)
@@ 3.2.1 Self-Reported Pain
Patients consistently reported lower pain scores when more people were in the room (Table 3). In table 3, pain groups were combined based on whether the pain scores were classified into the no pain and mild pain groups versus the moderate and severe pain groups. The Kruskal-Wallis H-test [kruskal_use_1952] was also used to compare the four pain groups separately, but no significant difference was found ($p=0.9$). The proportion of frames with the patient lying in bed was higher for the no and mild pain group compared to the moderate and severe pain group, but the difference wasn’t statistically significant.
@@
Model Metrics No/Mild pain Moderate/High Pain p-value
@@ 3.2.2 Patient Acuity
Model Metrics Stable Unstable p-value
@@ 3.2.3 Delirium
Model Metrics Non-Delirious Delirium p-value
@@ 4 Discussion
In this study, we employed the state-of-the-art YOLOv8 deep learning model [jocher_yolo_2023] to detect patient postures. The results demonstrated that the model had a harder time detecting the ”lying in bed class compared to the standing class. This result aligns with our expectations because individuals lying in bed were mostly covered by thick covers that made it difficult to distinguish which features belonged to the person.
@@
Among the metrics extracted from the model, hospital room visitations average and variance exhibited substantial correlations with measures of patient outcomes. These findings are logical because noise is a crucial factor associated with sleep deprivation and the onset of delirium [mistraletti_sleep_2008, cupka_effect_2020], and each individual represents a unique source of noise. This observation highlights the importance of using noninvasive autonomous systems to monitor patients. Individuals who tend to wake easily during sleep may be disturbed when healthcare providers open doors or inspect equipment [al_mutair_sleep_2020, bihari_factors_2012, eliassen_sleep_2011]. Minimizing the need for healthcare providers to enter or exit hospital rooms would allow patients to experience uninterrupted sleep and facilitate their recovery. Furthermore, implementing a noninvasive system for tracking patient mobility would allow care providers to remotely assess whether a patient is asleep and delay room entry until the patient is awake.
@@
Another intriguing discovery was the difference in the relationships identified when comparing subjective and objective measures of patient outcomes. Higher levels of hospital room visitations were linked to adverse objective measures of patient outcomes, including patient acuity and delirium. Conversely, self-reported pain exhibited a contrasting pattern. According to Boring et al., patients tend to downplay their pain when in the presence of friends or family members, as they wish to avoid causing undue worry [boring_how_2021]. Since we couldn’t identify anyone in each frame, it wasn’t possible to know whether the people in the room were care providers or family members. Incorporating each person’s relationship to the patient in video sequences would make it possible to explore the impact of family presence on self-reported pain records further.
@@
While a negative correlation was observed between patient outcomes and the number of people in each room, the current analysis does not provide adequate evidence to establish a causal relationship. The patient might have pressed the nurse call button [galinato_perspectives_2015] to request aid or their condition may have worsened to the extent that emergency medical staff were called in by the patient’s monitoring equipment. This would suggest that the number of individuals in the room resulted from the patients’ compromised outcomes. Incorporating the timing of each event that could potentially disturb a patient is needed to evaluate whether there is a causal relationship between hospital room visitations and patients’ outcomes.
@@
Using the average and variance of the number of people in an ICU room only considers one factor that can adversely affect patients’ sleep. Other stimuli, such as light or other sources of noise, have also been shown to impact patients’ sleeping habits [al_mutair_sleep_2020, bihari_factors_2012, eliassen_sleep_2011]. Incorporating real-time data from noise and light sensors would improve our system’s capacity to identify the critical factors contributing to sleep deprivation and notify care providers of potentially serious events, such as the onset of delirium.
@@
There are advantages and disadvantages associated with the utilization of depth cameras within an ICU as opposed to RGB cameras. Depth cameras provide 3D information about a scene and preserve the privacy of each person in a scene as long as the resolution is low. However, identifying which person is the patient can be a challenging endeavor when employing a depth camera. In the future, we will incorporate patient tracking using one of the object tracking architectures that have been integrated into the YOLOv8 model [zhang_bytetrack_2022, aharon_bot-sort_2022]. Using tracking information, we can follow the person lying in the hospital bed and evaluate their posture whenever they move around the hospital room. Knowing which person the patient is will enable better mobility characterization throughout the patient’s recovery period in the ICU. We can establish a comprehensive mobility index by amalgamating this data with information from other modalities, such as accelerometer data. This index would be valuable for preventing adverse events and enhancing patient outcomes.
@@ 5 Conclusions
6 Acknowledgements
A.B, P.R., and T.O.B. were supported by NIH/NINDS R01 NS120924, NIH/NIBIB R01 EB029699. P.R. was also supported by NSF CAREER 1750192.
[…]